OpenStudy (jpes2193):

Help with functions where you have F(X)= ## find what X is... Example in Comments

OpenStudy (jpes2193):

t(x)=0.25(x-7900)+79 The equation represents the tax bill [t] of a single person who's adjusted gross income (AGI) is x dollars between 7900 and 31600 inclusive. a. What is a single filer's tax bill if AGI is x=20,000 b. What is a single filer's AGI if the tax bill is 2540 I really need help with b.

jimthompson5910 (jim_thompson5910):

b. What is a single filer's AGI if the tax bill is 2540 so they give you the tax bill of t(x) = 2540 and they want you to use this to find x

jimthompson5910 (jim_thompson5910):

t(x)=0.25(x-7900)+79 2540=0.25(x-7900)+79 ... replace t(x) with 2540 2540-79=0.25(x-7900)+79-79 2461 = 0.25(x-7900) I'll let you finish

OpenStudy (jpes2193):

ugh.. i was wayyyyyy over thinking this... thanks that helps a ton.

OpenStudy (jpes2193):

17744=x

jimthompson5910 (jim_thompson5910):

you are correct, so x = 17,744 is the adjusted gross income (AGI) when the tax bill is t(x) = 2540

OpenStudy (jpes2193):

okay so if I had to write this as an ordered pair it would be (17,744, 2450)

jimthompson5910 (jim_thompson5910):

sorry it didn't notify me til now

jimthompson5910 (jim_thompson5910):

yes but one important thing if you are going to write the numbers in an ordered pair, do NOT use a comma in 17,744. The fact that a commas is used to separate x and y means you cannot use one to make "17,744" more readable. (17,744, 2450) looks like an ordered triple, ie a point in 3d space so write (17744, 2450) instead